# Safe and Optimal N-Spacecraft Swarm Reconfiguration in Non-Keplerian Cislunar Orbits

### v1

@@ 4.1 Passive Safety via Drift Trajectories From Controlled Nodes
One way to achieve passive safety is to enforce the exclusion of the drift trajectory from the keep-out zone for a predefined period [29, 31, 32]. For example, the passive safety of the $j$-th deputy and the chief is formulated as
$\max_{m=0,...,N_{\text{safe}}}\sqrt{{\boldsymbol{\rho}_{km}^{(j)}}^{\top}P\boldsymbol{\rho}_{km}^{(j)}}-{\boldsymbol{\rho}_{km}^{(j)}}^{\top}P\boldsymbol{\rho}_{km}^{(j)}\leq 0,\quad k=0,...,N$ (19a)
where $\boldsymbol{\rho}_{km}^{(j)}=\prod_{l=0}^{m-1}\Phi(t_{l+1},t_{l})T_{k}\boldsymbol{\zeta}_{\text{ns},k}^{(j)},$ (19b)
where Eq. (19b) defines the natural propagation of the deputy after each controlled epoch, and Eq. (19a) ensures the exclusion of the drift trajectory from the KOZ at each discretized time step. Note that this is an equivalent formulation to the well-known quadratic form of exclusion constraint $\sqrt{{\boldsymbol{\rho}^{(j)}}^{\top}P\boldsymbol{\rho}^{(j)}}\geq 1$. In this paper, the KOZ is modeled as an ellipsoid. Its geometric property is characterized by the matrix $P=\text{diag}([1/x_{e}^{2},1/y_{e}^{2},1/z_{e}^{2},0,0,0])$, where $(x_{e},y_{e},z_{e})$ are the semi-axes of the ellipsoid in the TNW frame. To prevent the quadratic growth of the number of constraints along the drift period, the maximum operator is applied over the interval $m=0,...,N_{\text{safe}}$. Note that an instantaneous collision avoidance is achieved by considering only $m=0$. Since Eq. (19a) is nonconvex, the following convexification is performed with respect to the reference trajectory $\bar{\boldsymbol{\rho}}_{k}$ as follows [44, 31]:
$-\boldsymbol{a}^{\top}_{k}T_{k}\boldsymbol{\zeta}^{(j)}_{\text{ns},k}+b_{k}\leq 0,\quad k=0,...,N$ (20a)
where $\boldsymbol{a}^{\top}_{k}={\bar{\boldsymbol{\rho}}}_{k}^{{(j)}^{\top}}\prod_{l=0}^{m^{*}_{k}-1}\Phi(t_{l+1},t_{l})D^{\top}PD\left(\prod_{l=0}^{j^{*}_{k}-1}\Phi(t_{l+1},t_{l})\right)^{\top},\quad b_{k}=\sqrt{\boldsymbol{a}^{\top}_{k}\bar{\boldsymbol{\rho}}_{k}^{(j)}},\quad D=[\boldsymbol{I}_{3},\boldsymbol{0}_{3}],$ (20b)
$m^{*}_{k}=\operatorname*{argmax}_{m=0,...,N_{\text{safe}}}\sqrt{{\bar{\boldsymbol{\rho}}}_{km}^{{(j)}^{\top}}P\bar{\boldsymbol{\rho}}^{(j)}_{km}}-\bar{\boldsymbol{\rho}}_{km}^{{(j)}^{\top}}P\bar{\boldsymbol{\rho}}_{km}^{(j)},$ (20c)
@@ 4.2 Passive Safety via First-Order condition of QPRO
This paper proposes a novel set of constraints that satisfy passive safety during a transfer by individually enforcing the first-order condition of remaining on a QPRO. This approach ensures that the deputy always adheres to the surface of a QPRIT of arbitrary size (defined by $\epsilon$) throughout transfers between QRPOs, thereby achieving passive safety. The constraint is expressed as
$\|h_{k}^{(j)}\|\leq\epsilon_{h},\quad\|\dot{h}_{k}^{(j)}\|\leq\epsilon_{\dot{h}},\quad\|\dot{\alpha}_{k}^{(j)}\|\leq\epsilon_{\dot{\alpha}},\quad\|\dot{\beta}_{k}^{(j)}\|\leq\epsilon_{\dot{\beta}},\quad k=0,...,N,\forall j$ (21a)
@@
The deputy’s condition to be on a QPRO is $h=\dot{\alpha}=\dot{\beta}=\dot{h}=0$. However, with the strict form of this equality, the problem would be infeasible as LTC becomes uncontrollable (cf. Eq. (13)). Instead, a set of $l_{2}$-norm constraints at each time step is added as shown in Eq. (21a). Furthermore, Eq. (21b) ensures that the deputy is outside of the QPRIT’s surface that corresponds to the terminal QPRO. To alleviate nonconvexity, the following convexification using the reference state $\bar{\alpha}_{k}$ and $\bar{\beta}_{k}$ is introduced:
@@
In this proposed approach, safety between agents is achieved solely via passively safe QPRO designs. Therefore, the above multi-agent OCP could be solved in a decentralized fashion, once the boundary conditions of each agent are defined. If there exits a strict requirement for the minimum separation between agents, a relative orbit design problem must be conducted by iteratively solving the above problem. However, the placement of safe relative orbits can be readily done by placing agents in the $\alpha-\beta$ plane of the local toroidal frame, similarly to the spacecraft swarm placement in ROE space [36], significantly simplifying the orbit design.
@@ 5 Results and Analysis
Table 2 summarizes the scenarios and parameters used for the respective trajectory optimization, such as reference orbits, boundary conditions, and constraint configurations. This paper adopts SCvx* [45] as an SCP algorithm to solve a nonconvex optimization. Note that because of the usage of a direct method with fixed time discretization, the strict fuel optimality [13, 46] is not attainable. However, as shown in the results, the transfer costs of reconfiguration in this study are on the order of $10^{0}$$10^{1}$ mm/s. Therefore, using $N=30$ for time discretization is justified, given that the station-keeping cost of the absolute orbit would be significantly higher. All numerical computation is performed using Intel Core i7-1260P.
@@
Orbits ($L_{2}$ South NRHO) 9:2 synodic (Fig. 10(a)) 4:1 sidereal (Fig. 10(b)) 3:1 synodic (Fig. 10(c)) 9:2 synodic (Fig. 10(a)) (Fig. 10(a))
@@
$t_{N}$ 2 revs. 2 revs. 2 revs.
$N$ 30 30 30
$[k_{p}^{-},k_{p}^{+}]$ [6,7,8,9], [21,22,23,24] [6,7,8,9], [21,22,23,24] [6,7,8,9], [21,22,23,24]
@@
QPRIT $\epsilon_{h}$ 1 m 1 m 1 m
$\epsilon_{\dot{\alpha}},\epsilon_{\dot{\beta}},\epsilon_{\dot{h}}$ 50 mm/s 50 mm/s 50 mm/s
DRIFT $T_{\text{safe}}$ 1 rev. MPC $Q$ position: 5e-8 m/s velocity: 5e-10 m/s2
@@
In this case study, if there are multiple oscillatory modes in a reference (chief) orbit, one with a smaller maximum along-track separation of the QPRIT is chosen to construct LTC. Such a torus structure is more desirable because a highly eccentric invariant circle requires a larger torus to ensure the minimum separation around the chief.
@@ 5.1 Comparison of Passively Safe Reconfiguration Strategies in the CR3BP
(b) TN plane projection
(c) NW plane projection
(d) TW plane projection
Figure 9: The terminal QPRIT (orange) and the keep-out zone (green) in the TNW frame. Although the planar separation is not guaranteed, the terminal QPRIT is passively safe from the defined keep-out zone.
@@
Nonlinear error (terminal position) [m] 0.008 0.0167 0.280 0.501
@@
(b) TNW frame
(c) Control History (TNW frame)
@@
(b) TNW frame
(c) Control History (TNW frame)
@@
The results of the optimization are summarized in Table 3, which enumerates (i) upfront computation time of monodromy matrix, history of eigenvectors, and STMs, (ii) optimization runtime, (iii) one-revolution passive safety, (iv) fuel cost, and (v) nonlinear error of the optimized trajectory evaluated at its terminal position. Note that the computation of STMs and other parameters is shared between CVX, CVX-QPRIT, and SCP-QPRIT, whereas additional computation is required for the SCP-DRIFT to propagate STMs for drift trajectories. Both SCPs converged to the local optimum, and the optimized trajectories are illustrated in Figures 11 and 12, respectively. Trajectories are resolved in both the TNW frame and the LVLH frame, and the impulsive control profile is also shown in each figure. The sky blue and orange surfaces are the initial and terminal QPRIT, respectively. Bold ellipses attached to the QPRITs illustrate the invariant circle at the initial and terminal epoch. Furthermore, the control profile is propagated along the nonlinear relative dynamics of the CR3BP, and the resulting nonlinear trajectories (solid magenta) are compared to the optimized linear trajectories (dotted blue). Additionally, the passive safety constraint in Eq. (19a) is evaluated for each optimized trajectory and summarized in Fig. 13. The state below the red line indicates the satisfaction of the safety constraint. It is observed from the figures that the one-orbit passive safety is achieved in both optimal trajectories. The SCP-DRIFT converges to a total cost of $\Delta v_{tot}=5.220$ mm/s, whereas the SCP-QPRIT yields $\Delta v_{tot}=8.510$ mm/s. The transfer cost is higher in the SCP-QPRIT due to the conservative nature of the passive safety constraint. It is evident from Fig. 11(b) that the controlled trajectory is inside the surface of the QPRIT that corresponds to the terminal state and makes a tangent contact with the surface of the KOZ. In contrast, the converged solution of SCP-QPRIT ensures that the deputy not only remains outside of the terminal QPRIT but also leverages the dynamical structure by staying on the surface of QPRIT throughout the transfer. Thus, the worst (maximum) constraint value shown in Fig. 13(b) is below the constraint violation threshold and achieves one-revolution passive safety without explicitly constraining the future state. A comparison of Figures 11(c) and 12(c) highlights the distinct optimal control strategies. While the SCP-DRIFT yields a control history resembling a two-impulse solution, the SCP-QPRIT solution spreads small impulses across the transfer, ensuring adherence to the QPRIT surface at each time step and the smooth transition along the QPRIT surface from the initial to the terminal state. Identifying which thrust profiles are more prone to nonlinear errors is not straightforward. SCP-QPRIT applies thrusts near the perilune, where system dynamics are highly sensitive, while SCP-DRIFT uses a small number of larger impulses that can potentially violate linearization assumptions. However, the nonlinear errors at the terminal positions are overall negligible in the case study as shown in Table 3, as the total $\Delta v$ values are typically small. Given that the evaluation is purely open-loop, the results show that the linearized dynamics provide a sufficiently accurate approximation of the nonlinear dynamics for the problems considered.
@@ 5.2.1 ER3BP $L_{2}$ South 4:1 Halo Orbit (sidereal resonance)
In ER3BP, absolute periodic orbits only permit sidereal resonance due to the eccentricity of the Earth-Moon orbit. In this case study, the $L_{2}$ South 4:1 NRHO is chosen as the reference chief orbit, as shown in Fig. 10(b). The optimized passively-safe solution (SCPP-QPRIT) is presented in Fig. 15, where Fig. 15(a) is the trajectory shown in the TNW frame, Fig. 15(b) is the control history. Furthermore, Figs. 15(c) and 15(d) are the trajectories plotted in the LTC. Similarly to the previous subsection, Table 4 summarizes the optimization results in CVX, CVX-QPRIT, and SCP-QPRIT, respectively.
(a) TNW frame
(b) Control History (TNW frame)
@@
Nonlinear error (terminal position) [m] 0.038 0.666 0.0164 0.870 0.667 0.657
As shown in Fig. 15(a), the shape and orientation of the osculating invariant circles at the initial and terminal epoch are different because of the 4:1 resonance of the reference, while the transfer time is 2 revolutions. Still, the SCP-QPRIT trajectory successfully performs safe reconfiguration using the quasi-periodic structure, compared to the CVX solution that has a total cost of $\Delta v_{tot}=8.19$ mm/s. Particularly, the solution of CVX-QPRIT results in the trajectory that stays on QPRIT surfaces for the entire transfer, although the deputy intrudes inside the terminal QPRIT, as shown in its LTC history (cf. Fig. 15(c)) and also smaller minimum separation as shown in the TNW frame (cf. Fig. 15(a)). The total costs of CVX-QPRIT and SCP-QPRIT are $\Delta v_{tot}=13.33$ mm/s and 13.77 mm/s, respectively, where passive safety is achieved via a marginal increase in transfer cost. When propagated under the nonlinear ER3BP dynamics, the terminal positional error between the linearized trajectory and the open-loop nonlinear trajectory is 0.0164 m, validating the accuracy of the linearized relative dynamics model in the ER3BP.

### v2

@@ 4.1 Passive Safety via Drift Trajectories From Controlled Nodes
One way to achieve passive safety is to enforce the exclusion of the drift trajectory from the keep-out zone for a predefined period [47, 49, 50]. The geometric configuration of an ellipsoidal KOZ is characterized by the diagonal matrix $P$. Using this matrix, the exclusion from the KOZ is expressed as a quadratic inequality $\sqrt{{\boldsymbol{x}}^{\top}P\boldsymbol{x}}\geq 1$. In this paper, the KOZ exclusion constrained is defined within the VNB frame, where the ellipsoid matrix is defined as $\text{diag}([1/x_{e}^{2},1/y_{e}^{2},1/z_{e}^{2},0,0,0])$ and $(x_{e},y_{e},z_{e})$ are the semi-axes of the ellipsoid in the VNB frame. Therefore, the passive safety of the $j$-th deputy and the chief is rewritten as [57]
$\max_{m=0,...,N_{\text{safe}}}\sqrt{{\boldsymbol{x}_{km}^{(j)}}^{\top}P\boldsymbol{x}_{km}^{(j)}}-{\boldsymbol{x}_{km}^{(j)}}^{\top}P\boldsymbol{x}_{km}^{(j)}\leq 0,\quad k=0,...,N,$ (22a)
where $\boldsymbol{x}_{km}^{(j)}=\prod_{l=0}^{m-1}\Phi(t_{l+1},t_{l})T_{k}\boldsymbol{\zeta}_{\text{ns},k}^{(j)},$ (22b)
where Eq. (22b) defines the natural propagation of the deputy after each controlled epoch, and Eq. (22a) ensures the exclusion of the drift trajectory from the KOZ at each discretized time step. In this subsection, for simplicity of notation, $\boldsymbol{x}$ represents the matrix representation of a state vector within the local frame. Note that this is an equivalent formulation to the well-known quadratic form of the exclusion constraint To prevent the quadratic growth of the number of constraints along the drift period, the maximum operator is applied over the interval $m=0,...,N_{\text{safe}}$. Note that an instantaneous collision avoidance is achieved by considering only $m=0$. Since Eq. (22a) is nonconvex, the following convexification is performed with respect to the reference orbit resolved in a local frame $\bar{\boldsymbol{x}}_{k}$ as follows [57, 49]:
$-\boldsymbol{a}^{\top}_{k}T_{k}\boldsymbol{\zeta}^{(j)}_{\text{ns},k}+b_{k}\leq 0,\qquad k=0,...,N$ (23a)
where $\boldsymbol{a}^{\top}_{k}={\bar{\boldsymbol{x}}}_{k}^{{(j)}^{\top}}\prod_{l=0}^{m^{*}_{k}-1}\Phi(t_{l+1},t_{l})D^{\top}PD\left(\prod_{l=0}^{j^{*}_{k}-1}\Phi(t_{l+1},t_{l})\right)^{\top},\qquad b_{k}=\sqrt{\boldsymbol{a}^{\top}_{k}\bar{\boldsymbol{x}}_{k}^{(j)}},\qquad D=[\boldsymbol{I}_{3},\boldsymbol{0}_{3}],$ (23b)
$m^{*}_{k}=\operatorname*{argmax}_{m=0,...,N_{\text{safe}}}\sqrt{{\bar{\boldsymbol{x}}}_{km}^{{(j)}^{\top}}P\bar{\boldsymbol{x}}^{(j)}_{km}}-\bar{\boldsymbol{x}}_{km}^{{(j)}^{\top}}P\bar{\boldsymbol{x}}_{km}^{(j)},$ (23c)
@@ 4.2 Passive Safety via First-Order condition of QPRO
A new constraint formulation is proposed in this paper to guarantee passive safety throughout the transfer, achieved by applying the first-order condition for staying on a QPRO. This approach ensures that the deputy always adheres to the surface of a QPRIT of arbitrary size (defined by $\varepsilon$) throughout transfers between QRPOs, thereby achieving passive safety. The constraint is expressed as
$\|h_{k}^{(j)}\|\leq\Delta_{h},\quad\|\dot{\alpha}_{k}^{(j)}\|\leq\Delta_{\dot{\alpha}},\quad\|\dot{\beta}_{k}^{(j)}\|\leq\Delta_{\dot{\beta}},\quad\|\dot{h}_{k}^{(j)}\|\leq\Delta_{\dot{h}},\quad k=0,...,N,\forall j$ (24a)
@@
The deputy’s condition to be on a QPRO is $h=\dot{\alpha}=\dot{\beta}=\dot{h}=0$. However, with the strict form of this equality, the problem would be infeasible as LTC becomes stationary (cf. Eq. (14)). Therefore, this constriant is relaxed to $l_{2}$-norm inequality at each time step as shown in Eq. (24a), where $\{\Delta_{h},\Delta_{\dot{\alpha}},\Delta_{\dot{\beta}},\Delta_{\dot{h}}\}$ are user-defined small parameters. The choice of these slack parameters depends on the length of the passive safety period and the accuracy of the dynamics modeling. Furthermore, Eq. (24b) ensures that the deputy is outside of the QPRIT’s surface that corresponds to the terminal QPRO. To alleviate nonconvexity, the following convexification using the reference state $\bar{\alpha}_{k}$ and $\bar{\beta}_{k}$ is introduced:
@@
In this proposed approach, safety between agents is achieved solely via passively safe QPRO designs. Therefore, the above multi-agent OCP could be solved in a decentralized fashion, once the boundary conditions of each agent are defined. If there exists a strict requirement for the minimum separation between agents, a relative orbit design problem must be conducted by iteratively solving the above problem. However, the placement of safe relative orbits can be readily done by placing agents in the $\alpha-\beta$ plane of the local toroidal frame, similarly to the spacecraft swarm placement in ROE space [54], significantly simplifying the QRPO design.
@@ 5 Results and Analysis
Table 2 summarizes the scenarios and parameters used for the respective trajectory optimization, such as reference orbits, boundary conditions, and constraint configurations. This paper adopts SCvx* [58] as an SCP algorithm to solve a nonconvex optimization. Note that because of the usage of a direct method with fixed time discretization, the strict fuel optimality [16, 59] is not attainable. However, as shown in the results, the transfer costs of reconfiguration in this study are on the order of $10^{0}$$10^{1}$ mm/s. Therefore, using $N=30$ for time discretization is justified, given that the station-keeping cost of the reference orbit would be significantly higher [60]. All numerical computation is performed using Intel Core i7-1260P.
@@
Orbits
($L_{2}$ South NRHO)
9:2 synodic
(Fig. 10(a))
4:1 sidereal
(Fig. 10(b))
3:1 synodic
(Fig. 10(c))
9:2 synodic
(Fig. 10(a))
(Fig. 10(a))
@@
$t_{f}$ 2 revs. 2 revs. 2 revs.
NN 30 30 30
$k_{p}^{-},k_{p}^{+}] [$6,9], [21,24] [6,9], [21,24] [6,9], [21,24]
@@
QPRIT $\Delta_{h}$ 1 m 1 m 1 m
$\Delta_{\dot{\alpha}},\Delta_{\dot{\beta}},\Delta_{\dot{h}}$ 50 mm/s 50 mm/s 50 mm/s
DRIFT $T_{\text{safe}}$ 1 rev. MPC $Q$
position: 5e-8 m/s
velocity: 5e-10 m/s2
@@
In this case study, if there are multiple oscillatory modes in a reference orbit, one with a smaller maximum along-track separation of the QPRIT is chosen to construct LTC. Such a torus structure is more desirable because a highly eccentric invariant circle requires a larger torus to ensure the minimum separation around the chief.
@@ 5.1 Comparison of Passively Safe Reconfiguration Strategies in the CR3BP
(b) VN plane projection
(c) NB plane projection
(d) BV plane projection
Figure 9: Keep-out zone (green) in the VNB frame and the passively-safe terminal QPRIT (orange).
@@
Nonlinear error
(terminal position) [m]
0.008 0.0167 0.280 0.501
@@
(b) VNB frame
(c) Control history (VNB frame)
@@
(b) VNB
frame
(c) Control History (VNB
frame)
@@
The results of the optimization are summarized in Table 3, which enumerates (i) upfront computation time of monodromy matrix, history of eigenvectors, and STMs, (ii) optimization runtime, (iii) one-revolution passive safety, (iv) fuel cost, and (v) nonlinear error of the optimized trajectory evaluated at its terminal position. Note that the computation of STMs and other parameters is shared between CVX, CVX-QPRIT, and SCP-QPRIT, whereas additional computation is required for the SCP-DRIFT to propagate STMs for drift trajectories. Both SCPs converged to the local optimum, and the optimized trajectories are illustrated in Figures 11 and 12, respectively. Trajectories are resolved in both the VNB frame and the LVLH frame, and the impulsive control profile is also shown in each figure. The sky blue and orange surfaces are the initial and terminal QPRIT, respectively. Bold ellipses attached to the QPRITs illustrate the invariant circle at the initial and terminal epoch. Furthermore, the control profile is propagated along the nonlinear relative dynamics of the CR3BP, and the resulting nonlinear trajectories (solid magenta) are compared to the optimized linear trajectories (dotted blue). Additionally, the passive safety constraint in Eq. (22a) is evaluated for each optimized trajectory and summarized in Fig. 13. The state below the red line indicates the satisfaction of the safety constraint. It is observed from the figures that the one-orbit passive safety is achieved in both optimal trajectories. The SCP-DRIFT converges to a total cost of $\Delta v_{tot}=5.220$ mm/s, whereas the SCP-QPRIT yields $\Delta v_{tot}=8.510$ mm/s. The transfer cost is higher in the SCP-QPRIT due to the conservative nature of the passive safety constraint. It is evident from Fig. 11(b) that the controlled trajectory is inside the surface of the QPRIT that corresponds to the terminal state and makes a tangent contact with the surface of the KOZ. In contrast, the converged solution of SCP-QPRIT ensures that the deputy not only remains outside of the terminal QPRIT but also leverages the dynamical structure by staying on the surface of QPRIT throughout the transfer. Thus, the worst (maximum) constraint value shown in Fig. 13(b) is below the constraint violation threshold and achieves one-revolution passive safety without explicitly constraining the future state. A comparison of Figures 11(c) and 12(c) highlights the distinct optimal control strategies. While the SCP-DRIFT yields a control history resembling a two-impulse solution, the SCP-QPRIT solution spreads small impulses across the transfer, ensuring adherence to the QPRIT surface at each time step and the smooth transition along the QPRIT surface from the initial to the terminal state. Identifying which thrust profiles are more prone to nonlinear errors is not straightforward. SCP-QPRIT applies thrusts near the perilune immediately before and after the no-control windows, where system dynamics are highly sensitive, while SCP-DRIFT uses a small number of larger impulses that can potentially violate linearization assumptions. However, the nonlinear errors at the terminal positions are overall negligible in the case study as shown in Table 3, as the total $\Delta v$ values are typically small. Given that the evaluation is purely open-loop, the results show that the linearized dynamics provide a sufficiently accurate approximation of the nonlinear dynamics for the problems considered.
@@ 5.2.1 ER3BP $L_{2}$ South 4:1 Halo Orbit (sidereal resonance)
In ER3BP, periodic orbits only permit sidereal resonance due to the eccentricity of the Earth-Moon orbit. In this case study, the $L_{2}$ South 4:1 NRHO is chosen as the reference orbit, as shown in Fig. 10(b). The optimized passively-safe solution (SCPP-QPRIT) is presented in Fig. 15, where Fig. 15(a) is the trajectory shown in the VNB frame, Fig. 15(b) is the control history. Furthermore, Figs. 15(c) and 15(d) are the trajectories plotted in the LTC. Similarly to the previous subsection, Table 4 summarizes the optimization results in CVX, CVX-QPRIT, and SCP-QPRIT, respectively.
(a) VNB
frame
(b) Control History (VNB
frame)
@@
Nonlinear error
(terminal position) [m]
0.038 0.666 0.0164 0.870 0.667 0.657
As shown in Fig. 15(a), the shape and orientation of the osculating invariant circles at the initial and terminal epoch are different because of the 4:1 resonance of the reference, while the transfer time is 2 revolutions. Still, the SCP-QPRIT trajectory successfully performs safe reconfiguration using the quasi-periodic structure, compared to the CVX solution that has a total cost of $\Delta v_{tot}=8.19$ mm/s. Particularly, the solution of CVX-QPRIT results in the trajectory that stays on QPRIT surfaces for the entire transfer, although the deputy intrudes inside the terminal QPRIT, as shown in its LTC history (cf. Fig. 15(c)) and also smaller minimum separation as shown in the VNB frame (cf. Fig. 15(a)). The total costs of CVX-QPRIT and SCP-QPRIT are $\Delta v_{tot}=13.33$ mm/s and 13.77 mm/s, respectively, where passive safety is achieved via a marginal increase in transfer cost. When propagated under the nonlinear ER3BP dynamics, the terminal positional error between the linearized trajectory and the open-loop nonlinear trajectory is 0.0164 m, validating the accuracy of the linearized relative dynamics model in the ER3BP.